OpenStudy (anonymous):

I am going to buy exotic fruits. Dragonfruit costs x-4 dollars. Starfruit is five dollars less expensive than rambutan. Rambutan costs 2x dollars more than dragonfruit. How much does it cost to buy one rambutan, two starfruit, and three dragonfruit? Your answer will be an expression that depends on x.

OpenStudy (anonymous):

how did you figure it out

OpenStudy (anonymous):

I made really bad typos

OpenStudy (anonymous):

Dragonfruit = x - 4 Starfruit = (x - 4) + 2x -5 Rambutan = (x - 4) + 2x

OpenStudy (anonymous):

oh never mind

OpenStudy (anonymous):

This probably won't make any sense but

OpenStudy (anonymous):

You pretty much just have to add it so so it looks like one big equation and then simplify (x - 4) + 2x + (x - 4) + 2x - 5 + (x - 4) + 2x - 5 + (x - 4) + (x - 4) + (x - 4)

OpenStudy (anonymous):

So (x - 4) + (2x) + (x - 4) + (2x) - (5) + (x - 4) + (2x) - (5) + (x - 4) + (x - 4) + (x - 4) @cj49 am I doing this right

OpenStudy (anonymous):

Well it's 'How much does it cost to buy one rambutan, two starfruit, and three dragonfruit?'

OpenStudy (anonymous):

HOW ABOUT THIS QUESTION GUYS please explain Juan has c cheesecakes. Clarence has 5 less than 3 times as many cheesecakes as Juan has and Darlene has 2 more than half as many cheesecakes as Clarence has. Find an expression in terms of c for the total number of cheesecakes Juan, Clarence, and Darlene have.

OpenStudy (cj49):

7x-17 x=17/7

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

Did you figure out the answer for the fruit one?

OpenStudy (anonymous):

oh yeah I got it sorry if it confused you

OpenStudy (anonymous):

It's okay I'm glad you got it! Good job